how many particles are present in 0.886 liters of unknown gas at standard temperature and pressure?
1
Expert's answer
2020-12-21T03:52:01-0500

N =( V/VM) * NA= (0.886/22.4)*6.02*1023=2.38*1022;

Answer: 2.38*1023 particles